OpenStudy (anonymous):

can someone help me understand why when factoring x(squared)-2x-80 the answer is (x+10)(x-8) but cannot be (x-10)(x+8)

OpenStudy (hoblos):

(x+10)(x-8) = x^2 -8x+10x-80 = x^2 +2x -80 (x-10)(x+8)= x^2 +8x-10x-80 = x^2 -2x -80 so it is (x-10)(x+8)

OpenStudy (jlastino):

The middle term should be the sum of the two numbers in each factor. The last term should be the product of the two...the correct answer should be (x-10) (x+8) because -10+8 = -2 and -10x8=-80